Anyone willing to tackle these two:
Please explain process. Bless you.
circumference = 240 inches; length of arc AB = 40 inches.Find the measure of the central angle ACB.
and
measurement of angle ACB = 30 degrees; length of arc AB = 100 millimeters, find the circumference.

Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(69443)

circumference = 240 inches; length of arc AB = 40 inches.Find the measure of the central angle ACB.
------------
The arc length = 1/6 of C, so the angle = (1/6)*360
= 60 degs
------------
and
measurement of angle ACB = 30 degrees; length of arc AB = 100 millimeters, find the circumference.
Angle = 30/360 = 1/12 of the circle.
100 mm = 1/12 of the circumference.
C = 1200 mm

Answer by lwsshak3(11628)

circumference = 240 inches; length of arc AB = 40 inches.
Find the measure of the central angle ACB.
..
circumference = 240=2πr
r=radius=240/2π=120/π
s=rx, s=length of arc, r=radius, x=central angle in radians
x=s/r=40/(120/π)=40π/120=π/3
central angle ACB=π/3 radians
..
measurement of angle ACB = 30 degrees; length of arc AB = 100 millimeters.
Find the circumference.
30º=π/6 radians
r=s/x=100/(π/6)=600/π
circumference=2πr=2π*600/π=1200 millimeters